If a rectangle measures 48 inches by 24 inches, what is the exact length in inches, of the diagonal of the rectangle?

Answer by rfer(16322)

c^2=48^2+24^2
c^2=2304+576
c^2=2880
c=sqrt 2880
c=53.66563146
I put all the decimals in because it wants exact length.